# Violation of Bell inequalities in an analogue black hole

### v1

@@ III Bipartite nonlocality
$\begin{split}\mbox{Tr}_{(0)}\Big{(}|0_{\omega=0}\rangle^{\rm in}\;{}^{\rm in}\langle 0_{\omega=0}|\Big{)}&=\\ \sum_{m,n,\mu,\nu}{\cal C}^{(0)}_{m,n,\mu,\nu}&\Big{(}|x^{-}_{m}x^{-}_{n}\rangle\,\langle x^{-}_{\mu}x^{-}_{\nu}|\\ &+|x^{+}_{m}x^{+}_{n}\rangle\,\langle x^{+}_{\mu}x^{+}_{\nu}|\Big{)}.\end{split}$ (30)
@@
If the eigenvalues $+1$ and $-1$ of operator $\hat{\Pi}_{x}$ were non-degenerate, this would impose $m=\mu$ and $n=\nu$ in expression (30) and the corresponding reduced state would be clearly separable. Nothing similar occurs in our situation, and indeed the reduced states are typically entangled, as shown in Sec. III.
@@
^$\begin{split}\hat{\mathscr{M}}^{(0|1|2)}(\omega)=&-\boldsymbol{a}\cdot\hat{\boldsymbol{\Pi}}^{(0)}\otimes\boldsymbol{b}\cdot\hat{\boldsymbol{\Pi}}^{(1)}\otimes\boldsymbol{c}\cdot\hat{\boldsymbol{\Pi}}^{(2)}\\ &+\boldsymbol{a}\cdot\hat{\boldsymbol{\Pi}}^{(0)}\otimes\boldsymbol{b}^{\prime}\cdot\hat{\boldsymbol{\Pi}}^{(1)}\otimes\boldsymbol{c}^{\prime}\cdot\hat{\boldsymbol{\Pi}}^{(2)}\\ &+\boldsymbol{a}^{\prime}\cdot\hat{\boldsymbol{\Pi}}^{(0)}\otimes\boldsymbol{b}\cdot\hat{\boldsymbol{\Pi}}^{(1)}\otimes\boldsymbol{c}^{\prime}\cdot\hat{\boldsymbol{\Pi}}^{(2)}\\ &+\boldsymbol{a}^{\prime}\cdot\hat{\boldsymbol{\Pi}}^{(0)}\otimes\boldsymbol{b}^{\prime}\cdot\hat{\boldsymbol{\Pi}}^{(1)}\otimes\boldsymbol{c}\cdot\hat{\boldsymbol{\Pi}}^{(2)}.\end{split}$ (33)
We note here that $\hat{\mathscr{S}}^{(0|1|2)}=\tfrac{1}{2}\hat{\mathscr{M}}^{(0|1|2)}+\tfrac{1}{2}\hat{\mathscr{M}}^{\prime}{}^{(0|1|2)}$, where ^$\hat{\mathscr{M}}^{\prime}{}^{(0|1|2)}$ is the same as ^$\hat{\mathscr{M}}^{(0|1|2)}$ with the prime reversed [76].
@@ V Conclusion
An important ingredient of our theoretical treatment is the assumption of Gaussianity which, in our case, results from the Bogoliubov approach. This is certainly a sensible first order hypothesis, but recent studies [83, 84, 85, 86] showed that nonlinear effects might significantly affect quantum emission processes in the context of analogue physics. It would then be of great interest to evaluate in detail the amplitude and the relevance of nonlinear back reaction effects on acoustic Hawking radiation in the black hole analogues we consider.
@@ Appendix A Different black hole configurations
The background classical field $\Phi$ of Eq. (2) is solution of a Gross-Pitaekskii equation which is the stationary and classical version of Eq. (1):
$-\frac{\hbar^{2}}{2m}\partial_{x}^{2}\Phi+[U(x)+g\,n-\mu]\Phi=0,$ (35)
where $n=|\Phi|^{2}$. An analogue black hole horizon is realized if the flow is subsonic far upstream (for $x\to-\infty$ in our convention) and takes the form of a downstream supersonic plane wave (for $x>0$). The corresponding classical field behaves as
@@
The second configuration is denoted as “waterfall”: The potential $U(x)$ is a step function and the upstream profile is half a dark soliton. It is close to the experimental realization of the Technion group [18, 19] and catches important aspects of the density-correlation pattern reported in [18], with nonetheless some caveats, see [52].
The third and last configuration is denoted as “delta peak”: the potential $U(x)$ is a repulsive Dirac distribution located at $x=0$ and the upstream profile is a part of a dark soliton, but not exactly one half of it, as is the case in the waterfall configuration.
@@ Appendix B Covariance matrices
[…]
@@
It has been argued in [20] that an efficient measure of bipartite entanglement was given by the “PPT measure” $\Lambda^{(i|j)}\equiv 1-\nu^{-}_{(i|j)}$ where $\nu^{-}_{(i|j)}$ is the lowest simplectic eigenvalue of the partial transpose of $\sigma^{(i|j)}$  [^5] The partial transposition corresponds to a mirror reflection in phase space which inverts the $p_{j}$ coordinate, leaving $q_{i}$, $p_{i}$ and $q_{j}$ unchanged [93]. The resulting covariance matrix can be brought by means of a symplectic transform to a diagonal form [94]. The corresponding diagonal elements are the simplectic eigenvalues. They are twice degenerate, and in our $4\times 4$ case there are thus two such eigenvalues: $\nu^{-}_{(i|j)}$ and $\nu^{+}_{(i|j)}$, with $\nu^{\pm}_{(i|j)}\in\mathbb{R}^{+}$ and $\nu^{-}_{(i|j)}\leq\nu^{+}_{(ij)}$.. The largest entanglement corresponds to $\Lambda^{(i|j)}=1$ while separability is reached when $\Lambda^{(i|j)}<0$. This separability condition can be shown to be equivalent to the Peres-Horodecki criterion [95, 96]. Contrarily to other observables, such as the Cauchy-Schwarz criterion or the generalized Peres-Horodecki parameter which have been often used in the domain, the PPT measure has the advantage of being an entanglement monotone. Other observables have been used in the context of analog gravity, which are monotonous measures of entanglement, such as the entanglement entropy, the entanglement of formation or the logarithmic negativity, but they all have some drawbacks: the state our our system is mixed in a finite temperature situation, which discards the entanglement entropy as a possible measure. The entanglement entropy generalizes for mixed states to the entanglement of formation [97], but this quantity is not easily determined in non symmetric two-mode Gaussian states such as the ones we consider [^6] In the situations we consider the reduced state of modes $i$ and $j$ is nonsymmetric, since in general the mixedness $a_{i}$ and $a_{j}$ (defined in (80)) are not equal.. The logarithmic negativity shares with the entanglement of formation the drawback of possibly violating monogamy inequalities [98]. This is a relatively mild drawback in the context of evaluating bipartite entanglement, but becomes prohibitive in the tripartite context. The Gaussian contangle was introduced in Ref. [98] as a quantity which has none of the previous deficiencies. It has been studied in the gravitational context [99] and also in analog gravity [20], but it is not of very practical use as its determination requires numerical minimisation of a complex expression. The PPT measure we use in the present study was introduced in Ref. [20] as a measure which mimics many aspects of the Gaussian contangle but is much simpler to evaluate. Its evaluation is as simple as the one of the above quoted quantities albeit it shares none of their drawbacks. The explicit expression of $\Lambda^{(i|2)}(\omega)$ characterizing the coupling between mode $i=0$ or $1$ and mode $j=2$ is given in Eq. (15) of the main text.
@@
$\begin{split}\langle\hat{c}^{\dagger}_{i}\hat{c}_{i}\rangle=&|S_{i0}|^{2}\bar{n}_{0}+|S_{i1}|^{2}\bar{n}_{1}+|S_{i2}|^{2}(1+\bar{n}_{2}),\\ \langle\hat{c}_{2}^{\dagger}\hat{c}_{2}\rangle=&|S_{20}|^{2}\bar{n}_{0}+|S_{21}|^{2}\bar{n}_{1}+|S_{22}|^{2}(1+\bar{n}_{2})-1,\\ \langle\hat{c}_{0}\hat{c}_{1}^{\dagger}\rangle=&S_{00}S_{10}^{*}\bar{n}_{0}+S_{01}S_{11}^{*}\bar{n}_{1}+S_{02}S_{12}^{*}(1+\bar{n}_{2}),\\ \langle\hat{c}_{i}\hat{c}_{2}\rangle=&S_{i0}S_{20}^{*}\bar{n}_{0}+S_{i1}S_{21}^{*}\bar{n}_{1}+S_{i2}S_{22}^{*}(1+\bar{n}_{2}),\\ \end{split}$ (46)
@@ Appendix C An analogue optical system
$\begin{split}&\langle\hat{f}_{0}^{\dagger}\hat{f}_{0}\rangle=\sum_{i=0}^{1}|S_{12}S_{0i}-S_{02}S_{1i}|^{2}\,\bar{n}_{i}/\sinh^{2}r_{2},\\ &\langle\hat{f}_{1}^{\dagger}\hat{f}_{1}\rangle=\cosh^{2}r_{2}\,\bar{n}_{01}+\sinh^{2}r_{2}(1+\bar{n}_{2}),\\ &\langle\hat{f}_{2}^{\dagger}\hat{f}_{2}\rangle=\sinh^{2}r_{2}\,\bar{n}_{01}+\cosh^{2}r_{2}(1+\bar{n}_{2})-1,\\ &\langle\hat{f}_{1}\hat{f}_{2}\rangle=\cosh r_{2}\sinh r_{2}(\bar{n}_{01}+\bar{n}_{2}+1).\\ \end{split}$ (56)
@@
Then, noticing that at zero temperature $\langle\hat{\Pi}^{(f_{1})}_{z}\otimes\hat{\Pi}^{(f_{2})}_{z}\rangle=1$ at all energies and that the vector $\boldsymbol{b}$, $\boldsymbol{b}^{\prime}$, $\boldsymbol{c}$ and $\boldsymbol{c}^{\prime}$ are not constrained with respect to each other, one can take these vectors to be $+\boldsymbol{e}_{z}$, thus obtaining
@@ Appendix D Eigenstates of the pseudo-spin operators
the eigenstates of $\hat{\Pi}_{z}$ are the number states, they have eigenvalue ±$\pm 1$ depending on their parity. In the following we will denote them as
@@
Contrarily to usual spins 1/2, the two eigenvalues (here $+1$ and $-1$) are infinitely degenerate. This property is shared by any projection of the pseudo-spin operator; it is in particular true for the operator $\hat{\Pi}_{x}$. We denote the eigenstates of $\hat{\Pi}_{x}$ associated to the eigenvalue ±$\pm 1$ by $|x^{\pm}_{n}\rangle$, they can be constructed by rotating the eigenstates $|z^{\pm}_{n}\rangle$ of $\hat{\Pi}_{z}$ by angles ±$\pm\pi/2$ around the $y$ axis:
@@
where ^$\hat{\cal R}_{y}(\theta)=\exp(-\frac{i}{2}\theta\,\hat{\Pi}_{y})$ is the operator of rotation of angle $\theta$ around the $y$ axis. The fact that $|x^{\pm}_{n}\rangle$ is an eigenstate of $\hat{\Pi}_{x}$ with eigenvalue ±$\pm 1$ is easily checked by direct application of $\hat{\Pi}_{x}$ to the left of expression (71) and use of the relations
@@ Appendix E Computing expectation values of observables
The use of the technique of Wigner transform is particularly well suited for determining the different contributions of the pseudo-spin operator to the CHSH measures (19) and (21). The reason is twofold: first, we consider a Gaussian state and for computing the required averages we thus just need to evaluate Gaussian integrals weighted by the Wigner transforms of the spin operators and secondly, the Wigner transforms of the pseudo-spins (16) are simple enough that the relevant integrals can be evaluated analytically.
@@
where ${\cal P}$ denotes the principal value. It is shown in the main text that the eigenvalues of the projections along a given axis of the pseudo-spin operator $\hat{\boldsymbol{\Pi}}$ are ±$\pm 1$. The above expressions of the Wigner transforms thus demonstrate that, contrarily to $\hat{\Pi}_{x}$, operators $\hat{\Pi}_{y}$ and $\hat{\Pi}_{z}$ are “improper” in the sense of Ref. [105]: they may be involved in violation of Bell inequality even for a state with a non-negative Wigner transform, such as the Gaussian state we consider.
@@
$\begin{split}\langle\hat{\Pi}^{(i)}_{x}\otimes\hat{\Pi}^{(2)}_{x}\rangle=&\frac{2}{\pi}\arctan\frac{2\,\mbox{Re}\,\langle\hat{c}_{i}\hat{c}_{2}\rangle}{\sqrt{a_{i}a_{2}-4(\mbox{Re}\,\langle\hat{c}_{i}\hat{c}_{2}\rangle)^{2}}},\\ \langle\hat{\Pi}^{(i)}_{y}\otimes\hat{\Pi}^{(2)}_{y}\rangle=&\frac{-1}{A_{i2}}\langle\hat{\Pi}^{(i)}_{x}\otimes\hat{\Pi}^{(2)}_{x}\rangle,\\ \langle\hat{\Pi}^{(i)}_{x}\otimes\hat{\Pi}^{(2)}_{y}\rangle=&\frac{2}{\pi\,a_{2}}\operatorname{arsinh}\frac{2\,\mbox{Im}\,\langle\hat{c}_{i}\hat{c}_{2}\rangle}{\sqrt{A_{i2}}},\\ \langle\hat{\Pi}^{(i)}_{y}\otimes\hat{\Pi}^{(2)}_{x}\rangle=&\frac{2}{\pi\,a_{i}}\operatorname{arsinh}\frac{2\,\mbox{Im}\,\langle\hat{c}_{i}\hat{c}_{2}\rangle}{\sqrt{A_{i2}}},\\ \end{split}$ (81)
@@
Appendix F Analytic maximization of ^$\langle\hat{\mathscr{B}}{}^{(i|2)}\rangle$
@@
$\begin{split}{\cal T}_{xx}=\;&\frac{2}{\pi}\arctan\frac{2\,|\langle\hat{c}_{i}\hat{c}_{2}\rangle|}{\sqrt{A_{i2}}}=-A_{i2}{\cal T}_{yy},\\ {\cal T}_{xy}=\;&{\cal T}_{yx}=0,\\ \end{split}$ (100)
@@
Consequently, we resort to a numerical method to evaluate the maximal violation of Bell inequalities and determine the corresponding optimal orientations for the vectors $\boldsymbol{a}$, $\boldsymbol{a}^{\prime}$, $\boldsymbol{b}$, $\boldsymbol{b}^{\prime}$, $\boldsymbol{c}$, and $\boldsymbol{c}^{\prime}$. More explicitly, we use a genetic algorithm which has proved very efficient for optimizing a function over a large parameter space [109]. This algorithm is based on natural selection: the code starts with a random set of solutions (in our case each of them consists of 12 parameters) which form all together what we call a population. We then compute, for each set of vectors ($\boldsymbol{a},\boldsymbol{a}^{\prime},\boldsymbol{b},\boldsymbol{b}^{\prime},\boldsymbol{c},\boldsymbol{c}^{\prime}$), the expectation value $\langle\hat{\mathscr{S}}{}^{(0|1|2)}\rangle$ by means of the technique exposed in Appendix E. For instance the contribution of the term $\langle\boldsymbol{a}\cdot\hat{\boldsymbol{\Pi}}{}^{(0)}\otimes\boldsymbol{b}\cdot\hat{\boldsymbol{\Pi}}{}^{(1)}\otimes\boldsymbol{c}\cdot\hat{\boldsymbol{\Pi}}{}^{(2)}\rangle$ to ^$\langle\hat{\mathscr{B}}{}^{(0|1|2)}\rangle$ can be evaluated from the knowledge of the terms ${\cal T}_{rst}$ defined in Eq. (82):
@@
[…]
@@
with similar formulae for the quantities $\hat{B}$, $\hat{B}^{\prime}$ and $\hat{C}$, $\hat{C}^{\prime}$. Since the operators $\hat{A}$, $\hat{B}$ and $\hat{C}$ operate in different Hilbert spaces, one can assume without loss of generality that all the vector products of type (105a) appearing in (104) are colinear with $\boldsymbol{e}_{z}$. In this case all the unit vectors $\boldsymbol{a}$, $\boldsymbol{a}^{\prime}$, $\boldsymbol{b}$, $\boldsymbol{b}^{\prime}$, $\boldsymbol{c}$ and $\boldsymbol{c}^{\prime}$ lie in the $xy$ plane. Denoting as $\theta_{a}$ the angle between $\boldsymbol{a}^{\prime}$ and $\boldsymbol{a}$, $\theta_{b}$ the angle between $\boldsymbol{b}^{\prime}$ and $\boldsymbol{b}$, and $\theta_{c}$ the angle between $\boldsymbol{c}^{\prime}$ and $\boldsymbol{c}$ (all these angles being in $[0,\pi]$), one gets
[…]
It is a simple matter to check that when $\theta_{a}$, $\theta_{b}$ and $\theta_{c}$ run through $[0,2\pi]^{3}$ the eigenvalues of the operator appearing in the right hand side of the above formula are all comprised in $[0,8]$. The extremal values 0 and 8 are reached for $(\theta_{a},\theta_{b},\theta_{c})=(0,0,\pi)$ and $(\pi/2,\pi/2,\pi/2)$, respectively. It then follows that for any choice of the measurement directions $\boldsymbol{a}$, $\boldsymbol{a}^{\prime}$, $\boldsymbol{b}$, $\boldsymbol{b}^{\prime}$, $\boldsymbol{c}$ and $\boldsymbol{c}^{\prime}$
@@
[…]

### v2

@@ III Bipartite nonlocality
[…]
@@
If the eigenvalues +1+1 and 1-1 of operator $\hat{\Pi}_{x}$ were non-degenerate, this would impose $m=\mu$ and $n=\nu$ in expression (30) and the corresponding reduced state would be clearly separable. Nothing similar occurs in our situation, and indeed the reduced states are typically entangled, as shown in Sec. III.
@@
$\begin{split}\hat{\mathscr{M}}^{(0|1|2)}(\omega)=&-\boldsymbol{a}\cdot\hat{\boldsymbol{\Pi}}^{(0)}\otimes\boldsymbol{b}\cdot\hat{\boldsymbol{\Pi}}^{(1)}\otimes\boldsymbol{c}\cdot\hat{\boldsymbol{\Pi}}^{(2)}\\ &+\boldsymbol{a}\cdot\hat{\boldsymbol{\Pi}}^{(0)}\otimes\boldsymbol{b}^{\prime}\cdot\hat{\boldsymbol{\Pi}}^{(1)}\otimes\boldsymbol{c}^{\prime}\cdot\hat{\boldsymbol{\Pi}}^{(2)}\\ &+\boldsymbol{a}^{\prime}\cdot\hat{\boldsymbol{\Pi}}^{(0)}\otimes\boldsymbol{b}\cdot\hat{\boldsymbol{\Pi}}^{(1)}\otimes\boldsymbol{c}^{\prime}\cdot\hat{\boldsymbol{\Pi}}^{(2)}\\ &+\boldsymbol{a}^{\prime}\cdot\hat{\boldsymbol{\Pi}}^{(0)}\otimes\boldsymbol{b}^{\prime}\cdot\hat{\boldsymbol{\Pi}}^{(1)}\otimes\boldsymbol{c}\cdot\hat{\boldsymbol{\Pi}}^{(2)}.\end{split}$ (33)
We note here that $\hat{\mathscr{S}}^{(0|1|2)}=\tfrac{1}{2}\hat{\mathscr{M}}^{(0|1|2)}+\tfrac{1}{2}\hat{\mathscr{M}}^{\prime}{}^{(0|1|2)}$, where $\hat{\mathscr{M}}^{\prime}{}^{(0|1|2)}$ is the same as $\hat{\mathscr{M}}^{(0|1|2)}$ with the prime reversed [79].
@@ V Conclusion
In an analog system, the essential ingredient for observing Hawking radiation is roughly the same as the one initially invoked by Hawking in the gravitational context [38]: the vacuum of the outgoing modes is not the same as the vacuum of the ingoing modes. In the BEC system we consider, this is physically due to the fact that the asymptotic upstream region is subsonic while the asymptotic downstream region is supersonic. A flow of this type induces the mismatch of the asymptotic dispersion relations illustrated in Fig. 1, with additional (negative norm) channels in the downstream region, resulting in the Bogoliubov transformation (5) from which stems relation (8) between the ingoing and outgoing vacua. The other important physical ingredient is the sonic character of the dispersion relation in the long wavelength limit. This results in the low energy divergence of the coefficients of the $S$ matrix involving the incoming negative norm mode (the $S_{j,2}$ coefficients with our notations) [86] and, in fine, in the finite Hawking temperature [39, 40]. This being said, the Bogoliubov transform we consider is not an exotic one; it is attached to a standard quadratic Bose Hamiltonian [35] and to the production of correlated pairs of quasi-particles. This is the reason why, as illustrated in Fig. 2, we can mimic our system by a simple optical setup involving a standard parametric downconvertion process. From this perspective the existence of three and not simply two modes (which might be considered as atypical in the Bogoliubov context) simply stems from the presence of a beam-splitter in the optical setting. Going back to the gravitational point of view, it has been argued in [22] that this beam splitter embodies the scattering of Hawking radiation by the geometry of the black hole, i.e., it effectively reproduces the greybody factor. The fact that such a simple pair production process induces genuine tripartite entanglement and genuine tripartite non-locality is far from being intuitive and is an important results of Ref. [22] and of the present study.
It results from our Bogoliubov treatment that we describe the system as a Gaussian state. This is certainly a sensible first order hypothesis, but recent studies [87, 88, 89, 90] showed that nonlinear effects might significantly affect quantum emission processes in the context of analogue physics. It would then be of great interest to evaluate in detail the amplitude and the relevance of nonlinear back reaction effects on acoustic Hawking radiation in the black hole analogues we consider.
@@ Appendix A Different black hole configurations
In this Appendix we first briefly present the range of validity of expansion (2) and then introduce some stationary solutions of the classical background field $\Phi(x)$ corresponding to an analog black hole.
For a quasi-1D guided BEC transversely trapped by a harmonic potential with angular frequency $\omega_{\perp}$, expansion (2) is valid in the “1D mean field regime” [31] defined by
$\left(\frac{a}{a_{\perp}}\right)^{2}\ll n_{\rm typ}a\ll 1,$ (35)
where $a$ is the 3D s-wave scattering length, $a_{\perp}=\sqrt{\hbar/m\omega_{\perp}}$ and $n_{\rm typ}$ is the typical linear density (number of atoms per unit length). The left inequality in (35) ensures that the finite phase coherence length (induced by long wavelength quantum fluctuations) is exponentially large compared to the healing length. The inequality at the riht ensures that the transverse degrees of freedom are frozen. For a transverse trap of frequency of 1 kHz, one gets $(a_{\perp}/a)^{2}=1.7\times 10^{-5}$ and $2.6\times 10^{-4}$, for 23Na and 87Rb, respectively. Hence the domain of validity of the 1D mean field approximation used in the present work typically ranges over four orders of magnitudes in density.
In the regime where (35) holds, the background classical field $\Phi$ of Eq. (2) is solution of a Gross-Pitaekskii equation which is the stationary and classical version of Eq. (1):
$\frac{\hbar^{2}}{2m}\partial_{x}^{2}\Phi+[U(x)+g\,n-\mu]\Phi=0,$ (36)
where $n=|\Phi|^{2}$ and $g=2\hbar\omega_{\perp}a$. An analogue black hole horizon is realized if the flow is subsonic far upstream (for $x\to$-$\infty$ in our convention) and takes the form of a downstream supersonic plane wave (for $x>0$). The corresponding classical field behaves as
@@
The second configuration is denoted as “waterfall”: The potential $U(x)$ is a step function and the upstream profile $\phi_{u}(x)$ is half that of a dark soliton. It is close to the experimental realization of the Technion group [20, 21] and catches important aspects of the density-correlation pattern reported in [20], with nonetheless some caveats, see [55].
The third and last configuration is denoted as “delta peak”: the potential $U(x)$ is a repulsive Dirac distribution located at $x=0$ and the upstream profile $\phi_{u}(x)$ is a part of a dark soliton, but not exactly one half of it, as is the case in the waterfall configuration.
@@ Appendix B Covariance matrices
$\sigma_{i}=\left(1+2\,\langle\hat{c}_{i}^{\dagger}\hat{c}_{i}\rangle\right)\,\mathds{1}_{2},$ (42)
@@
It has been argued in [22] that an efficient measure of bipartite entanglement was given by the “PPT measure” $\Lambda^{(i|j)}\equiv 1-\nu^{-}_{(i|j)}$ where $\nu^{-}_{(i|j)}$ is the lowest simplectic eigenvalue of the partial transpose of $\sigma^{(i|j)}$  [^5] The partial transposition corresponds to a mirror reflection in phase space which inverts the $p_{j}$ coordinate, leaving $q_{i}$, $p_{i}$ and $q_{j}$ unchanged [97]. The resulting covariance matrix can be brought by means of a symplectic transform to a diagonal form [98]. The corresponding diagonal elements are the simplectic eigenvalues. They are twice degenerate, and in our $4\times 4$ case there are thus two such eigenvalues: $\nu^{-}_{(i|j)}$ and $\nu^{+}_{(i|j)}$, with $\nu^{\pm}_{(i|j)}\in\mathbb{R}^{+}$ and $\nu^{-}_{(i|j)}\leq\nu^{+}_{(ij)}$.. The largest entanglement corresponds to $\Lambda^{(i|j)}=1$ while separability is reached when $\Lambda^{(i|j)}<0$. This separability condition can be shown to be equivalent to the Peres-Horodecki criterion [99, 100]. Contrarily to other observables, such as the Cauchy-Schwarz criterion or the generalized Peres-Horodecki parameter which have been often used in the domain, the PPT measure has the advantage of being an entanglement monotone. Other observables have been used in the context of analog gravity, which are monotonous measures of entanglement, such as the entanglement entropy, the entanglement of formation or the logarithmic negativity, but they all have some drawbacks: the state our our system is mixed in a finite temperature situation, which discards the entanglement entropy as a possible measure. The entanglement entropy generalizes for mixed states to the entanglement of formation [101], but this quantity is not easily determined in non symmetric two-mode Gaussian states such as the ones we consider [^6] In the situations we consider the reduced state of modes $i$ and $j$ is nonsymmetric, since in general the mixedness $a_{i}$ and $a_{j}$ (defined in (81)) are not equal.. The logarithmic negativity shares with the entanglement of formation the drawback of possibly violating monogamy inequalities [102]. This is a relatively mild drawback in the context of evaluating bipartite entanglement, but becomes prohibitive in the tripartite context. The Gaussian contangle was introduced in Ref. [102] as a quantity which has none of the previous deficiencies. It has been studied in the gravitational context [103] and also in analog gravity [22], but it is not of very practical use as its determination requires numerical minimisation of a complex expression. The PPT measure we use in the present study was introduced in Ref. [22] as a measure which mimics many aspects of the Gaussian contangle but is much simpler to evaluate. Its evaluation is as simple as the one of the above quoted quantities albeit it shares none of their drawbacks. The explicit expression of $\Lambda^{(i|2)}(\omega)$ characterizing the coupling between mode $i=0$ or 11 and mode $j=2$ is given in Eq. (15) of the main text.
@@
[…]
@@ Appendix C An analogue optical system
[…]
@@
Then, noticing that at zero temperature $\langle\hat{\Pi}^{(f_{1})}_{z}\otimes\hat{\Pi}^{(f_{2})}_{z}\rangle=1$ at all energies and that the vector $\boldsymbol{b}$, $\boldsymbol{b}^{\prime}$, $\boldsymbol{c}$ and $\boldsymbol{c}^{\prime}$ are not constrained with respect to each other, one can take these vectors to be $\boldsymbol{e}_{z}$, thus obtaining
@@ Appendix D Eigenstates of the pseudo-spin operators
the eigenstates of $\hat{\Pi}_{z}$ are the number states, they have eigenvalue $\pm 1$ depending on their parity. In the following we will denote them as
@@
Contrarily to usual spins 1/2, the two eigenvalues (here +1+1 and 1-1) are infinitely degenerate. This property is shared by any projection of the pseudo-spin operator; it is in particular true for the operator $\hat{\Pi}_{x}$. We denote the eigenstates of $\hat{\Pi}_{x}$ associated to the eigenvalue $\pm 1$ by $|x^{\pm}_{n}\rangle$, they can be constructed by rotating the eigenstates $|z^{\pm}_{n}\rangle$ of $\hat{\Pi}_{z}$ by angles $\pm\pi/2$ around the $y$ axis:
@@
where $\hat{\cal R}_{y}(\theta)=\exp(-\frac{i}{2}\theta\,\hat{\Pi}_{y})$ is the operator of rotation of angle $\theta$ around the $y$ axis. The fact that $|x^{\pm}_{n}\rangle$ is an eigenstate of $\hat{\Pi}_{x}$ with eigenvalue $\pm 1$ is easily checked by direct application of $\hat{\Pi}_{x}$ to the left of expression (72) and use of the relations
@@ Appendix E Computing expectation values of observables
The use of the technique of Wigner transform (see, e.g., Refs. [105, 106]) is particularly well suited for determining the different contributions of the pseudo-spin operator to the CHSH measures (19) and (21). The reason is twofold: first, we consider a Gaussian state and for computing the required averages we thus just need to evaluate Gaussian integrals weighted by the Wigner transforms of the spin operators and secondly, the Wigner transforms of the pseudo-spins (16) are simple enough that the relevant integrals can be evaluated analytically.
@@
where $\cal P$} denotes the principal value. It is shown in the main text that the eigenvalues of the projections along a given axis of the pseudo-spin operator $\hat{\boldsymbol{\Pi}}$ are $\pm 1$. The above expressions of the Wigner transforms thus demonstrate that, contrarily to $\hat{\Pi}_{x}$, operators $\hat{\Pi}_{y}$ and $\hat{\Pi}_{z}$ are “improper” in the sense of Ref. [111]: they may be involved in violation of Bell inequality even for a state with a non-negative Wigner transform, such as the Gaussian state we consider.
@@
[…]
@@
Appendix F Analytic maximization of $\langle\hat{\mathscr{B}}{}^{(i|2)}\rangle$
@@
[…]
@@
Consequently, we resort to a numerical method to evaluate the maximal violation of Bell inequalities and determine the corresponding optimal orientations for the vectors $\boldsymbol{a}$, $\boldsymbol{a}^{\prime}$, $\boldsymbol{b}$, $\boldsymbol{b}^{\prime}$, $\boldsymbol{c}$, and $\boldsymbol{c}^{\prime}$. More explicitly, we use a genetic algorithm which has proved very efficient for optimizing a function over a large parameter space [115]. This algorithm is based on natural selection: the code starts with a random set of solutions (in our case each of them consists of 12 parameters) which form all together what we call a population. We then compute, for each set of vectors ($\boldsymbol{a},\boldsymbol{a}^{\prime},\boldsymbol{b},\boldsymbol{b}^{\prime},\boldsymbol{c},\boldsymbol{c}^{\prime}$), the expectation value $\langle\hat{\mathscr{S}}{}^{(0|1|2)}\rangle$ by means of the technique exposed in Appendix E. For instance the contribution of the term $\langle\boldsymbol{a}\cdot\hat{\boldsymbol{\Pi}}{}^{(0)}\otimes\boldsymbol{b}\cdot\hat{\boldsymbol{\Pi}}{}^{(1)}\otimes\boldsymbol{c}\cdot\hat{\boldsymbol{\Pi}}{}^{(2)}\rangle$ to $\langle\hat{\mathscr{B}}{}^{(0|1|2)}\rangle$ can be evaluated from the knowledge of the terms ${\cal T}_{rst}$ defined in Eq. (83):
@@
where $\cdot$$\cdot$] and $\{\cdot,\cdot\}$ denote the commutator ans the anticommutator, respectively. From the SU(2) algebra of the pseudo-spins it is easily proven that
@@
with similar formulae for the quantities $\hat{B}$, $\hat{B}^{\prime}$ and $\hat{C}$, $\hat{C}^{\prime}$. Since the operators $\hat{A}$, $\hat{B}$ and $\hat{C}$ operate in different Hilbert spaces, one can assume without loss of generality that all the vector products of type (106a) appearing in (105) are colinear with $\boldsymbol{e}_{z}$. In this case all the unit vectors $\boldsymbol{a}$, $\boldsymbol{a}^{\prime}$, $\boldsymbol{b}$, $\boldsymbol{b}^{\prime}$, $\boldsymbol{c}$ and $\boldsymbol{c}^{\prime}$ lie in the $xy$ plane. Denoting as $\theta_{a}$ the angle between $\boldsymbol{a}^{\prime}$ and $\boldsymbol{a}$, $\theta_{b}$ the angle between $\boldsymbol{b}^{\prime}$ and $\boldsymbol{b}$, and $\theta_{c}$ the angle between $\boldsymbol{c}^{\prime}$ and $\boldsymbol{c}$ (all these angles being in $0,\pi]$), one gets
$\begin{split}\left(\hat{\mathscr{S}}{}^{(0|1|2)}\right)^{2}=&2+2\cos\theta_{a}\cos\theta_{b}\cos\theta_{c}\\ &+2\sin\theta_{a}\sin\theta_{b}\,\hat{\Pi}_{z}^{(0)}\otimes\hat{\Pi}_{z}^{(1)}\otimes\mathds{1}^{(2)}\\ &+2\sin\theta_{b}\sin\theta_{c}\,\mathds{1}^{(0)}\otimes\hat{\Pi}_{z}^{(1)}\otimes\hat{\Pi}_{z}^{(2)}\\ &+2\sin\theta_{a}\sin\theta_{c}\,\hat{\Pi}_{z}^{(0)}\otimes\mathds{1}^{(1)}\otimes\hat{\Pi}_{z}^{(2)}.\end{split}$ (107)
It is a simple matter to check that when $\theta_{a}$, $\theta_{b}$ and $\theta_{c}$ run through $0,2\pi]^{3}$ the eigenvalues of the operator appearing in the right hand side of the above formula are all comprised in $[0,8]$. The extremal values 0 and 8 are reached for $(\theta_{a},\theta_{b},\theta_{c})=(0,0,\pi)$ and $(\pi/2,\pi/2,\pi/2)$, respectively. It then follows that for any choice of the measurement directions $\boldsymbol{a}$, $\boldsymbol{a}^{\prime}$, $\boldsymbol{b}$, $\boldsymbol{b}^{\prime}$, $\boldsymbol{c}$ and $\boldsymbol{c}^{\prime}$
@@
$2\cos\theta\,\boldsymbol{a}_{+}=\boldsymbol{a}+\boldsymbol{a}^{\prime},\;\;2\sin\theta\,\boldsymbol{a}_{-}=\boldsymbol{a}-\boldsymbol{a}^{\prime},$ (110)